# KCL-PH-TH/2025-33

### v2

@@ 5 Summary and outlook
We have studied the evolution of scalar inhomogeneities during a kination epoch beyond the linear regime using numerical relativity. Starting with small perturbations, our results in the sub-horizon limit agree with perturbative analysis: linear perturbations grow relative to the background and eventually dominate the universe, but do not collapse into black holes. We then examined initially large perturbations and found that they rapidly begin to behave like radiation unless the initial overdensity is large $\delta_{0}\gg 1$, in which case gravitational forces are strong enough to induce immediate collapse into a black hole. However, such large sub-horizon perturbations cannot arise from the growth of initially small fluctuations after horizon re-entry: their domination over the background would halt further growth, preventing collapse. We therefore conclude that only mechanisms capable of sourcing large sub-horizon inhomogeneities at early times could lead to gravitational collapse.
On the other hand, super-horizon inhomogeneities manifest a much more interesting phenomenology with regard to strong gravity effects. The very long oscillation timescales of these modes imply that they behave like curvature perturbations prior to horizon re-entry. We observed that even in the limit where $\delta_{0}\ll 1$, such perturbations grow significantly relative to the background within just a few e-folds of kination. After horizon re-entry, their fate depends on their amplitude, they may either evolve into a radiation-like background–similar to typical for sub-horizon modes–or collapse into black holes. Our key finding is that the critical overdensity extrapolated from our numerical evolution, is much smaller compared to the perturbative predictions, although it converges to the perturbative predictions when the initial wavelength approaches $H_{0}^{-1}$. For increasingly larger wavelengths, we find that the extrapolated critical overdensity decreases, reaching around $\delta_{c,\text{numerical}}\approx 0.02$ at $\lambda_{0}\approx 2000H_{0}^{-1}$, as shown in Figure 4.
@@
In addition, we have chosen initial conditions corresponding to pure kination, where the scalar potential vanishes and no initial kinetic perturbations are present. While this provides a well-defined setup for studying scalar inhomogeneities, it would be valuable to explore scenarios in which a potential with an exponentially decaying region, as discussed in the Introduction, is included in conjunction with a dynamical geometry. Furthermore, it would be interesting to investigate perturbations in the kinetic sector with varying wavelengths, provided that the integrability conditions remain satisfied.

### v3

@@ 5 Summary and outlook
We have investigated the evolution of scalar inhomogeneities during a kination epoch beyond the linear regime using numerical relativity. Starting with small perturbations, our results in the sub-horizon limit agree with perturbative analysis: linear perturbations grow relative to the background and eventually dominate the universe, but do not collapse into black holes. We then examined initially large perturbations and found that they rapidly begin to behave like radiation unless the initial overdensity is large $\delta_{0}\gg 1$, in which case gravitational forces are strong enough to induce immediate collapse into a black hole. However, such large sub-horizon perturbations cannot arise from the growth of initially small fluctuations after horizon re-entry, since once they dominate over the background, further growth is suppressed, thereby preventing collapse. We therefore conclude that only mechanisms capable of sourcing large sub-horizon inhomogeneities at early times could lead to gravitational collapse.
On the other hand, super-horizon inhomogeneities exhibit a much richer phenomenology with respect to strong-gravity effects. The very long oscillation timescales of these modes imply that they behave like curvature perturbations prior to horizon re-entry. We observed that even in the limit where $\delta_{0}\ll 1$, such perturbations grow significantly relative to the background within just a few e-folds of kination. After horizon re-entry, their fate depends on their amplitude, they may either evolve into a radiation-like background–similar to typical sub-horizon modes–or collapse into black holes. Our key finding is that the critical overdensity extrapolated from our numerical evolution is much smaller than perturbative predictions, although it converges to the perturbative predictions when the initial wavelength approaches $H_{0}^{-1}$. For increasingly larger wavelengths, we find that the extrapolated critical overdensity decreases, reaching around $\delta_{c,\text{numerical}}\approx 0.02$ at $\lambda_{0}\approx 2000H_{0}^{-1}$, as shown in Figure 4.